# Phase transition like behaviors of Propagation of Passenger Stranding phenomena in Subway Networks

### 12 Jan 2024 v1

@@ 2.2.1 Dynamics Mechanism of Cluster Growth
Understanding the dynamic mechanism behind PPS generation is crucial for understanding the phase transition like behaviors in service systems and holds significant practical implications for enhancing the resilience of subway. The emergence of PPS in complex dynamical systems, comprising demand and supply basic units, is attributed to the complex and interconnected factors. In this study, the eigen microstate method of non-equilibrium statistical physics is used to uncover the characteristic dynamic patterns of PPS formation (see Method in the Appendix), which not only helps to understand the formation of clusters but also facilitates the study of cluster vanishing. include cluster formation and vanishing. According to the singular value decomposition (SVD)[39], the spatiotemporal correlation matrix A of stranded passenger number at each station can be factorized as $A=U\cdot\Sigma\cdot V^{T}$, where $U$ is the eigen microstate and $V$ is their evolution with time, $\Sigma$ is a diagonal matrix with the elements of $\sigma_{I}$. By equation ${{(A}_{I}^{e})}_{it}\equiv U_{I}\bigotimes V_{I}$, we get ${{(A}_{I}^{e})}_{it}$. We call the ensemble defined by ${{(A}_{I}^{e})}_{it}$ an eigen ensemble of system, where I represents the index of eigenstates, the subscript i denotes the station index, and t represents the time index. We consider $\sigma_{I}$ as the probability amplitude and $\omega_{I}=\sigma_{I}^{2}$ as the probability of the eigen ensemble ${{(A}_{I}^{e})}_{it}$ in the statistical ensemble A. The larger the probability $\omega_{I}$ associated with an eigen microstate, the greater its contribution to the formation of the PPS phenomenon. The eigen patterns with the largest eigenvalues are crucial factors determining the PPS phenomenon. By comparing the spatiotemporal characteristics of these key eigen microstates with the spatiotemporal characteristics of supply and demand, the micro mechanisms underlying the formation of these eigen patterns can be identified. Considering that the 4 four commuting corridors are independent in the spatial dimension, the eigen microstate method analyzing is applied separately to analyze Clusters I-VI.
@@
Taking Cluster I (Fig.6A) of the northern commuting corridor as an example (Fig.6A), this corridor serves two adjacent large residential centers, Huilongguan and Tiantongyuan, collectively known as the "Huitian area" for commuting demands. This area covers approximately 63 square kilometers and accommodates nearly 900,000 residents, earning the title of "Asia’s largest sleeping town." A large number of residents commute between the Huitian area and the city center by taking subway as their daily transportation. We concentrate on the morning rush hour, from 6:00 to 13:00 to reveal the formation and vanishing of Cluster I. We take a time snapshot every 10 minutes to record the number of stranded passengers at each station in Cluster I, resulting in a spatiotemporal correlation matrix of stranded passenger counts with 43 time snapshots. By performing SVD on this matrix, we obtain 8 independent eigen microstates ${{(A}_{I}^{e})}_{it}$ and their eigen values $\omega_{I}$. 8 $\omega_{I}$ are shown in Fig. 6B , and the sum of probabilities of the top three micro eigenstates reaches 99.983%. We consider these 3 patterns as the major mechanisms of PPS phenomenon which are named based on their potential causes as: Local Pattern, Nonlocal Pattern, and Topology Pattern, respectively. Similar analyses are conducted for Cluster II-IV, as well, leading to evolution mechanisms similar to Cluster I (see Supplement Meterials, B.4.1, Fig.B.18,B.19,B.20). It is worth noting that the eigen microstate method can be regarded as an unsupervised machine learning method that excavates hidden patterns from the Spatiotemporal data. The interpretations of these patterns often require a phenomenological and empirical perspective[40]. Therefore, while the eigen microstate method used in this study provides rigorous eigen modes, the interpretation of the underlying micro mechanisms and the naming of these patterns are somewhat subjective and serve as a reference.
@@
Figure 5: Station distribution and probabilities of eigen microstates for Cluster I. (A) Distribution of stations within Cluster I. The labeled stations include Station 200 (Huoying), Station 87 (LisuiQiao), Station 231 (BeiYuan), Station 232 (WangjingXi), Station 86 (Tiantong Yuan), Station 85 (Tiantong Yuan), Station 88 (Lishuiqiao Nan), and Station 89 (BeiyuanluBei). (B) Probabilities of the first eight eigen microstates corresponding to Cluster I during 6:00-13:00.
Local pattern: The largest eigenvalue, responsible for 84.88% of the total contribution, exhibits significant spatiotemporal variations at three distinct stations: i=87, 231, and 232. The temporal evolution of ${{(A}_{1}^{e})}_{it}$ at these stations is depicted in Figure 6A, demonstrating a characteristic pattern with a single peak during the early peak period, reaching its maximum around 8:50, and returning to the initial state by approximately 10:40. Station 231 exhibits the most prominent amplitude, while station 87 shows a more gradual variation. These observed patterns coincide with the temporal changes in the number of passengers entering the corresponding stations (i.e., demand scale), suggesting that this stranded pattern arises from an increase in the same passenger demand. Thus, we term this phenomenon the local pattern. Consequently, the primary driving factor behind the increase in stranded passengers at these stations is the mismatch between supply and demand caused by localized demand growth.
Figure 6: (A) Temporal evolution of the largest eigenstates in Cluster I for stations 87, 231, and 232 from 6:00 to 13:00. Stations 85, 86, 200, 88, and 89 do not exhibit significant variations in their eigenstates and are therefore not shown. (B) Variation in the number of entering passengers for stations 87, 231, and 232 from 6:00 to 13:00.
Nonlocal pattern: The second-largest eigenvalue accounts for 3.08% of the total variance. Figures 7(A), (C), and (E) depict the temporal evolution of ${{(A}_{2}^{e})}_{it}$ for stations 87, 231, and 232, respectively. In contrast to the characteristic single peak of Local pattern, this pattern exhibits alternating positive and negative variations, resembling the derivative of a single peak function. By comparing the morning rush hour stranded passenger counts at upstream stations i=86, 87, and 231 (Figures 7B, D, F) with ${{(A}_{2}^{e})}_{it}$, it can be inferred that the mechanism of stranded conditions involves propagation. This suggests that the variations in this pattern are driven by an increase in boarding passengers at upstream stations during rush hour periods. The stranded conditions at upstream stations result in a reduction of service supply downstream, leading to an increase in stranded passengers at downstream stations (Figure 1D). We named this phenomenon the Nonlocal pattern. Specifically, between 7:50 and 9:00, the increase in boarding passengers at station 86 leads to overcrowding, and this overcrowded condition does not improve at the nearest downstream station, 87. This interdependent relationship between adjacent stations facilitates the propagation of stranded conditions along the line direction. This effect is a secondary consequence of the Local pattern, resulting in a lower contribution from the Nonlocal pattern. Similarly, the increase in boarding passengers at station 231 exacerbates the stranded conditions downstream by reducing supply. During this period, both station 87 and station 231 experience stranded conditions, which have a suppressing effect on downstream station 232. Between 9:00 and 10:40, as the number of boarding passengers at station 231 increases, the stranded conditions at downstream station 232 worsen. Simultaneously, the decrease in boarding passengers at station 86 gradually reduces the occupation of supply at downstream stations, thereby alleviating the stranded conditions. Additionally, the impact of station 87 on the stranded conditions at station 231 starts to weaken. This mode reflects the propagation of stranded conditions and explains how the traffic supply at upstream stations influences the stranded conditions at downstream stations.
Figure 7: Variation of the second-largest eigenstates and passenger demand in Cluster I. (A), (C), and (E) depict the temporal evolution of the second-largest eigenstates for stations 87, 231, and 232, respectively, from 6:00 to 13:00. (B), (D), and (F) represent the changes in the number of boarding passengers at upstream stations 86, 87, and 231, respectively. The eigenstates of other stations within Cluster I exhibit minimal variations.
Topological pattern: The third-largest eigenvalue corresponds to a pattern contribution of 1.99%. Fig. 5 illustrates the temporal evolution of ${{(A}_{3}^{e})}_{it}$. This pattern unveils the influence of the infrastructure’s topological structure on passenger stranding, which we refer to as the Topological pattern. The distinguishing characteristic of this pattern is the significant variation observed at a specific station, designated as Local pattern, as depicted in Fig. 9 for station 87. Unlike the other 2 stations, station 87 functions as a transfer station, with multiple upstream stations contributing to the propagation of stranding and multiple downstream stations contributing to stranding alleviation, while stations 231 and 232 serve as non-transfer stations. Consistent with the findings from steady-state condition calculations under spatially and temporally homogeneous demand conditions discussed in the previous section, stations with high network betweenness centrality at transfer points play a pivotal role in PPS occurring. Under steady-state conditions, only the topological structure of the network influences PPS. In this system characterized by spatially and temporally heterogeneous supply and demand, the contribution of the topological factor ranks third.
2.2.2 Validating the Mechanism of Cluster Formation
The match between supply scale and demand scale is a fundamental factor determining the state of a system. To validate the interpretation of 3 pattern of cluster formation, addition simulations with varying the supply scale and fixing the demand scale are performed. The supply scale is controlled by varying the train frequencies on both Line 5 and Line 13, where Cluster I is located, while maintaining the spatiotemporal distribution characteristics of the supply. This allowed us to observe the inhibiting effect of increased supply scale on the both Local and Nonlocal patterns. Detailed information on the train departure frequency control scheme can be found in Supplement Meterials, B.4.2. By employing the eigen microstate method, the Local, Nonlocal, and Topological patterns of Cluster I under different service supply scales Ps (see Supplement Meterials, B.4.3, Fig.B.22) are analyzed. As shown in Fig. 9A, an increase in the supply scale leads to a continuous rise in the contribution of the Topological pattern, while the contributions of the Local and Nonlocal patterns are suppressed. In these simulations, the topological structure of the network remains unchanged, and any absolute changes in stranded passengers resulting from the topological structure can be neglected. The increase in the contribution of the topological pattern in Fig. 10A is only attributed to the decreased contributions of the local and nonlocal patterns. This supports our reasonable assumption that the local and nonlocal patterns arise due to a mismatch between supply scale and demand scale.
@@ 2.2.2 Validating the Mechanism of Cluster Formation
When $P_{s}>1.5$, the Nonlocal pattern, an important mechanism in PPS, is almost completely suppressed, and the contribution of the Local pattern is significantly reduced. Compared to the Nonlocal pattern, the local pattern is more sensitive to changes in supply scale. When $Ps\approx 1.3$, the contribution of the local pattern rapidly decreases, indicating that the increased supply scale directly fulfills the local supply shortage. During this process, the contribution of the Nonlocal pattern decreases slowly. The Local pattern is highly sensitive to supply improvements, hence an increase in demand leads to a decrease in the Local pattern. The triggering factor for the Nonlocal pattern is the presence of local stranding, which allows the propagation of the Nonlocal pattern through the PPS mechanism to downstream stations. Therefore, the regulation of the Nonlocal pattern by the supply scale is indirect, with the response to supply increases occurring later than that of the Local pattern. It is worth noting that when the supply scale exceeds 1.5, the contribution of the Nonlocal pattern is almost completely eliminated, and further improvements in the supply scale lead to slower enhancements in the Local pattern.
@@
The results validate the reasonableness of the cluster formation mechanism analyzed above and would provide guidance for decision making of policies. Increasing the supply scale within a certain range is beneficial for suppressing both the Local and Nonlocal patterns, and the influence on the Nonlocal pattern gradually strengthens as the supply increases between 1.3 and 1.5. After the supply scale exceeds 1.5, the inhibiting effect of the Local pattern significantly weakens, and the Nonlocal pattern is completely eliminated. Increasing the supply scale has limited effect on suppressing the Local and Nonlocal patterns. If further improvement in the system’s service efficiency is desired, it is necessary to shift the regulatory policies towards optimizing the network’s topological structure.
Figure 9: The proportion of local pattern (green), nonlocal pattern (red), and network topology pattern (blue) in the total variance as a function of $P_{s}$.
3 Conclusion
@@
In this study, we examine the subway system as a complex system that consists of 2 type of basic units, namely passengers and subway trains with a service relationship. By employing computational simulations of a big-data-driven model system, we explore the dynamics of stranding phenomenon within the subway system, particularly how it evolves in response to variations in the scale of supply. Our computational results unveil analogous behaviors of phase transitions in both service efficiency, as quantified by the average waiting time for passengers, and the scale of stranding phenomenon within the subway. We identify a transition point, which quantitatively characterizes the threshold of subway service resilience failure. Employing the method of eigen microstates from non-equilibrium statistical physics, we elucidate the formation mechanisms of PPS phenomenon. These achievements enhance the understanding of the features of service efficiency in subway networks and provides a theoretical foundation for strategies such as topological planning and optimization of train schedules to improve service. The proposed framework, involving complex systems comprising multiple interdependent basic units with service relationships, offers a valuable perspective for studying urban systems. In the pursuit of smart cities, this framework serves as a reference for evaluating urban planning, governmental policies, and business decisions. Furthermore, the methodology used in present work has broader implications for similar scheduling and logistics management challenges encountered in other service systems, including demand-driven or flexible bus scheduling, elevator dispatching in skyscrapers, supply chain management, medical appointment assignment, logistics management for aircraft carriers, and container logistics management.
Supplement Meterials

### 28 Feb 2024 v2

@@ 2.3 Dynamics mechanism of cluster growth
Understanding the dynamic mechanism behind PPS generation is crucial for understanding the phase transition like behaviors in service systems and holds significant practical implications for enhancing the resilience of the subway. The emergence of PPS in complex dynamical systems, comprising demand and supply basic units, is attributed to the complex and interconnected factors. In this study, the eigen microstate method of non-equilibrium statistical physics is used to uncover the characteristic dynamic patterns of PPS formation (see “Method”for details), which not only helps to understand the formation of clusters but also facilitates the study of cluster vanishing. include cluster formation and vanishing. According to the singular value decomposition (SVD)38, the spatiotemporal correlation matrix $\bm{A}$ of stranded passenger number at each station can be factorized as $\bm{A}=\bm{U}\cdot\bm{\Sigma}\cdot\bm{V}^{\mathrm{T}}$, where $\bm{U}$ is the eigen microstate and $\bm{V}$ is their evolution with time, $\bm{\Sigma}$ is a diagonal matrix with the elements of $\sigma_{I}$. By equation ${{\bm{A}}_{I}^{\mathrm{e}}}\equiv\bm{U_{I}}\bigotimes\bm{V_{I}}$, we get ${{\bm{A}}_{I}^{\mathrm{e}}}$. We call the ensemble defined by ${{\bm{A}}_{I}^{\mathrm{e}}}$ an eigen ensemble of system, where $I$ represents the index of eigenstates. We consider $\sigma_{I}$ as the probability amplitude and $\omega_{I}=\sigma_{I}^{2}$ as the probability of the eigen ensemble ${{\bm{A}}_{I}^{\mathrm{e}}}$ in the statistical ensemble $\bm{A}$. The larger the probability $\omega_{I}$ associated with an eigen microstate, the greater its contribution to the formation of the PPS phenomenon. The eigen patterns with the largest eigenvalues are crucial factors determining the PPS phenomenon. By comparing the spatiotemporal characteristics of these key eigen microstates with the spatiotemporal characteristics of supply and demand, the micro mechanisms underlying the formation of these eigen patterns can be identified. Considering that the 4 commuting corridors are independent in the spatial dimension, the eigen microstate method analyzing is applied separately to analyze Clusters I-VI.
@@
Taking Cluster I (Fig. 5a) of the northern commuting corridor as an example, this corridor serves 2 adjacent large residential centers, Huilongguan and Tiantongyuan, collectively known as the ”Huitian area” for commuting demands. This area covers approximately 63 square kilometers and accommodates nearly 900,000 residents, earning the title of ”Asia’s largest sleeping town”. A large number of residents commute between the Huitian area and the city center by taking subway as their daily transportation. We concentrate on the morning rush hour, from 6:00 to 13:00 to reveal the formation and vanishing of Cluster I. We take a time snapshot every 10 minutes to record the number of stranded passengers at each station in Cluster I, resulting in a spatiotemporal correlation matrix of stranded passenger counts with 43 time snapshots. By performing SVD on this matrix, we obtain 8 independent eigen microstates ${{(A}_{I}^{e})}_{it}$ and their eigen values $\omega_{I}$. 8 $\omega_{I}$ are shown in Fig. 5b , and the sum of probabilities of the top three micro eigenstates reaches 99.983%. We consider these 3 patterns, named as the Local Pattern, Nonlocal Pattern, and Topology Pattern based on their potential causes (more details in Supplementary Information, B.4.1), as the major mechanisms of the PPS phenomenon. They provide explanations for the impact of passenger demand on PPS, the propagation mechanism of PPS, and the impact of network topology on PPS. Similar analyses are conducted for Cluster II-IV, as well, leading to evolution mechanisms similar to Cluster I (see Supplementary Information, B.4.2, Fig. B.12 - Fig. B.14). It is worth noting that the eigen microstate method can be regarded as an unsupervised machine learning method that excavates hidden patterns from the Spatiotemporal data. The interpretations of these patterns often require a phenomenological and empirical perspective39. Therefore, while the eigen microstate method used in this study provides rigorous eigen modes, the interpretation of the underlying micro mechanisms and the naming of these patterns are somewhat subjective and serve as a reference.
@@
Figure 5: Station distribution and probabilities of eigen microstates for Cluster I. a Distribution of stations and critical stations within Cluster I. The labeled critical stations include Station 200 (Huoying), Station 87 (LisuiQiao), Station 231 (BeiYuan), Station 232 (WangjingXi), Station 86 (Tiantong Yuan), Station 85 (Tiantong Yuan), Station 88 (Lishuiqiao Nan), and Station 89 (BeiyuanluBei). b Probabilities of the first 8-th eigen microstates corresponding to Cluster I during 6:00-13:00.
2.4 The impact of supply and demand on cluster growth
The equilibrium between supply and demand is a fundamental factor that determines the state of the system. Therefore, we investigate the changes in the influence of different patterns by manipulating the demand scale and supply scale. In the present study, we maintain the spatiotemporal characteristics of the supply and the ratio of scheduled train departures to actual train departures is used as the system’s service supply scale, denoted as $P_{\mathrm{s}}$. The details of the train departure frequency regulation scheme can be found in the Supplementary Information, B.4.3. The eigen microstate method is used to analyze the Local, Nonlocal, and Topological patterns of Cluster I. Fig. 6a illustrates the proportional contribution of Topological pattern depending on both $P_{\mathrm{r}}$ and $P_{\mathrm{s}}$ which equals to the combined Local and Nonlocal patterns by considering the normalization condition of these 3 contributions. The blue area represents supply surplus, while the red area represents supply shortage. In the region where $P_{\mathrm{r}}$ is relatively small and $P_{\mathrm{s}}$ is relatively large, the system does not exhibit stranding due to excessive supply. Thus, eigen microstate method is not applicable, as indicated by the blank area in Fig. 6a. The black dashed line in the green region corresponds to the boundary of the combined changes in the Local and Nonlocal patterns. Below this boundary, the Topological pattern dominans the PPS phenomena, and the influence of the Local and Nonlocal patterns is minimal. With an increasing ratio of $P_{\mathrm{s}}$ over $P_{\mathrm{r}}$, the influence of the Local and Nonlocal patterns gradually amplifies, confirming that these 2 patterns arise from the mismatch between supply scale and demand scale.
Figure 6: a Contribution of Topological pattern depending on scales of supply and demand. b Contributions of Local pattern, Nonlocal pattern and Topological pattern in the total variance for different values of $P_{\mathrm{s}}$. c Contributions of Local pattern, Nonlocal pattern and Topological pattern in the total variance for different values of $P_{\mathrm{r}}$.
2.4.1 The impact of supply scale on patterns
To observe the impact of supply scale regulation on stranding phenomenon, we consider Cluster I as an example. Under the condition with constant demand scale, we increase the supply scale on Line 5 and Line 13, where Cluster I is located, to enhance the supply scale while maintaining the spatiotemporal distribution characteristics of the supply. This allows us to observe the inhibitory effect of increased supply scale on the Local and Nonlocal patterns caused by demand surplus (see Fig. 6b), corresponding to the white dashed line parallel to the x-axis in Fig. 6a at $P_{\mathrm{r}}=1.0$. As shown in Fig. 6b, with the increase in the supply scale, the contribution of the Topological pattern continues to rise, while both the Local and Nonlocal patterns are suppressed. In Fig. 6b, the increase in the contribution of the Topological pattern comes at the expense of decreased contributions from the Local and Nonlocal patterns. This indicates the validity of associating the Local and Nonlocal patterns with supply shortages.
@@ 2.4.1 The impact of supply scale on patterns
When $P_{\mathrm{s}}>1.5$, the Nonlocal pattern, which is an important mechanism in the PPS, is almost completely suppressed, and the contribution of the Local pattern is significantly reduced. Compared to the Nonlocal pattern, the Local pattern is more sensitive to changes in the supply scale. When $P_{\mathrm{s}}\approx 1.3$, the contribution of the Local pattern decreases rapidly, indicating that the increase in supply scale directly satisfies the supply. Meanwhile, the contribution of the Nonlocal pattern decreases slowly. When $P_{\mathrm{s}}\in[1.3,1.5]$, the Local and Nonlocal patterns exhibit mutual influences. With the increase of supply scale, the Nonlocal pattern shows a weak peak that initially increases and then decreases, while the Local pattern exhibits a valley. It is important to note that with the increase in supply, the absolute of stranding scale decreases, but the contributions of the Local and Nonlocal patterns slightly increase. The Local pattern is highly sensitive to supply change, leading to a decrease as demand increases. The trigger for the Nonlocal pattern is the stranding generated by the Local pattern, allowing it to propagate downstream through the PPS mechanism. Therefore, the regulation of the Nonlocal pattern by supply scale is indirect, with the response to an increase in supply lagging behind that of the Local pattern. It is worth noting that when supply exceeds 1.5, the contribution of the Nonlocal pattern is almost completely eliminated, while the improvement of the Local pattern becomes slower with further increases in the supply scale.
@@
The above results validate the rationality of the cluster growth mechanism analyzed earlier and provide valuable insights for regulatory policies. Increasing the supply within a certain range is advantageous for suppressing both the Local and Nonlocal patterns. The influence on the Nonlocal pattern gradually strengthens as the supply increases to the range of 1.3-1.5. However, once the supply exceeds 1.5, the inhibitory effect of the Local pattern significantly diminishes, and the Nonlocal pattern is completely eliminated. Increased transportation supply has limited impact on suppressing the Local and Nonlocal patterns. To further enhance the system’s service efficiency, it is recommended to shift the focus of regulatory policies towards optimizing the network’s topological structure.
2.4.2 The impact of demand scale on patterns
In the context of a constant supply, accommodating passenger demand is also an important issue. Therefore, we maintain the supply at its actual scale and regulate the demand scale $P_{\mathrm{r}}$, while observing the contributions of the 3 patterns. The results are presented in Fig. 6c which corresponds to the white dashed line parallel to the y-axis line in Fig. 6a representing $P_{\mathrm{s}}=1.0$. The demand regulation method aligns with the approach described in section 2.3.1. As illustrated in Fig. 6c, after $P_{\mathrm{r}}$ exceeds 0.6, an increase in $P_{\mathrm{r}}$ leads to the Local and Nonlocal patterns becoming the dominant factors contributing to passenger stranding. Both the Local and Nonlocal patterns are enhanced, while the influence of the Topological pattern diminishes significantly, with its contribution nearly eliminated. This further confirms that the Local and Nonlocal patterns originate from the matching between demand and supply. The results suggest that the supply scale is effective in suppressing the Local and Nonlocal patterns when the demand is high, particularly in suppressing the Local pattern. When the demand scale is less than 0.6, increasing the supply has limited impact on suppressing the Local and Nonlocal patterns; therefore, alternative approaches, such as modifying the network structure to enhance system efficiency, should be considered.
Furthermore, we observed that as $P_{\mathrm{r}}$ increases, the Topological pattern and the Local pattern of Cluster I respectively dominate at different stages, while the influence of the Local pattern in Clusters II-IV remains significant. This is due to the fact that Cluster I was constructed earlier and has a relatively complex topological structure. Stations with high demand and transfer stations are located in close proximity, resulting in the combined impact of multiple patterns on passenger stranding. Moreover, during the process of changing $P_{\mathrm{r}}$, the Topological pattern in this cluster has a relatively larger contribution compared to other clusters. On the other hand, the TOD pattern is more pronounced in Clusters II-IV, which consist of a larger number of non-transfer stations. Consequently, the Local pattern caused by demand remains prominent in these clusters (for the analysis of the evolution mechanism in Clusters II-IV, please refer to SI, 2.4.5, Fig. S13, Fig. S14).
3 Discussion
@@
In this study, we examine the subway system as a complex system that consists of 2 types of basic units, namely passengers and subway trains with a service relationship. By employing computational simulations of a big-data-driven model system, we explore the dynamics of stranding phenomenon within the subway system, particularly how it evolves in response to variations in the scale of supply. Our computational results unveil analogous behaviors of phase transitions in both service efficiency, as quantified by the average waiting time for passengers, and the scale of stranding phenomenon within the subway. We identify a transition point, which quantitatively characterizes the threshold of subway service resilience failure. Employing the method of eigen microstates from non-equilibrium statistical physics, we elucidate the formation mechanisms of the PPS phenomenon. These achievements enhance the understanding of the features of service efficiency in subway networks and provides a theoretical foundation for strategies such as topological planning and optimization of train schedules to improve service. The proposed framework, involving complex systems comprising multiple interdependent basic units with service relationships, offers a valuable perspective for studying urban systems. In the pursuit of smart cities, this framework serves as a reference for evaluating urban planning, governmental policies, and business decisions. Furthermore, the methodology used in the present work has broader implications for similar scheduling and logistics management challenges encountered in other service systems, including demand-driven or flexible bus scheduling, elevator dispatching in skyscrapers, supply chain management, medical appointment assignment, logistics management for aircraft carriers, and container logistics management.
4 Method
We begin by adopting Gibbs’ ensemble method 20 and define a statistical ensemble comprising $M$ microstates from a complex system with $N$ components. This statistical system is represented by a normalized $N\times M$ ensemble matrix $\bm{A}$, in which each row corresponds to the number of stranded passengers $S_{i}$($t$) ($i$ = 1, $M$) at different stations for a specific time snapshot. In this context, $M$ is set to 234, while the number of time samples is $N$ = 43.
At time $t$, the number of stranded passengers at station i is denoted as SPD, which represents the quantity of passengers. To calculate the average value of SPD at station $i$ over $M$ time periods, we compute the mean value as
$\langle S_{i}\rangle=\frac{1}{M}\sum_{i=1}^{M}S_{i}(t).$ (1)
At a certain time $t$, the agent $i$ has a fluctuation
$\delta S_{i}(t)=S_{i}(t)-\langle S_{i}\rangle.$ (2)
We define a microstate with fluctuations of all agents, which is represented by a normalized $N$-dimensional vector21
$\delta\bm{S}(t)=\begin{bmatrix}\delta S_{1}(t)\\ \delta S_{2}(t)\\ \vdots\\ \delta S_{N}(t)\end{bmatrix}.$ (3)
With the $M$ microstates, we can compose a statistical ensemble of the complex system. This ensemble is described by an $N\times M$ matrix $\bm{A}$ with elements
$A_{it}=\frac{\delta S_{i}(t)}{\sqrt{c_{0}}},$ (4)
where $C_{0}=\sum_{t=1}^{M}\sum_{i=1}^{N}\delta S_{i}^{2}(t)$. The column order of $\bm{A}$ is in accord with the evolution of microstate. As in Reference40, the correlation between the microstates at $t$ and $t^{\prime}$ is defined by their vector product
$C_{tt^{\prime}}=\delta S\left(t\right)^{\mathrm{T}}\cdot\delta S\left(t^{\prime}\right)=\sum_{i=1}^{N}{\delta S_{i}\left(t\right)\delta S_{i}\left(t^{\prime}\right)}.$ (5)
With $C_{tt^{\prime}}$ as its elements, we can get an M × M correlation matrix of microstate as
$\bm{C}=C_{0}\bm{A}^{\mathrm{T}}\cdot\bm{A},$ (6)
whose trace $\mathrm{Tr}C=\sum_{t=1}^{M}C_{tt}=C_{0}$.
According to the singular value decomposition (SVD) 41, the ensemble matrix $\bm{A}$ can be factorized as
$\bm{A}=\bm{U}\cdot\bm{\Sigma}\cdot\bm{V}^{\mathrm{T}},$ (7)
where $\delta$ is an $N\times M$ diagonal matrix with elements
$\delta_{ij}=\begin{cases}\delta_{I},&\text{for $i=j\leq r$}\\ 0,&\text{for otherwise},\end{cases},$ (8)
where $r=min\left(N,M\right)$.
We can rewrite the ensemble matrix $\bm{A}$ as
$\bm{A}=\sum_{I=1}^{r}{\sigma_{I}\bm{A_{I}}^{\mathrm{e}}},$ (9)
where $\bm{A_{I}}^{\mathrm{e}}\equiv\bm{U_{I}}\bigotimes\bm{V_{I}}$ is an $N$$\times$ $M$ matrix with elements ${{(A}_{I}^{\mathrm{e}})}_{it}=U_{it}V_{it}$. We call the ensemble defined by $A_{I}^{e}$ an eigen ensemble of system.
From ${\mathrm{Tr}}C=C_{0}$, we have the relation
$\sum_{I=1}^{r}\sigma_{I}^{2}=1.$ (10)
Therefore, we consider $\sigma_{I}$ as the probability amplitude and ${\omega_{I}}=\sigma_{I}^{2}$ as the probability of the eigen ensemble $\bm{A_{I}}^{\mathrm{e}}$ in the statistical ensemble $\bm{A}$. This is analogous to quantum mechanics, where a wave function can be written in terms of eigen functions. The square of the absolute value of expansion coefficient is the probability of the corresponding eigen microstate.
Acknowledgements
This work was supported by the Beijing Social Science Foundation (No. 21JCB093), Major Program of The Humanity and Social Science Fund of Beijing Jiaotong University (No.2023JBW1002) and National Key Research and Development Program of China (NKPs; Nos. 2021YFA1000300 and 2021YFA1000303).
Author contributions
Xinghua Zhang conceived the idea. Xinghua Zhang and Xinyi Li developed the methodology. Shengda Zhao implemented the computer code and supported the algorithms. Liang Wang applied statistical and mathematical techniques to analyze the study data. Qing Wang prepared the published work, specifically focusing on visualization. Liang Wang and Daqing Gong provided the data. Xun Zhang,Fang Liu and Xiaodong Zhang created the models. Xinghua Zhang and Xinyi Li wrote the paper. All authors have reviewed and approved the manuscript.
Competing interests
The authors declare no competing interests.